# Attacking the Spike: On the Transferability and Security of Spiking Neural Networks to Adversarial Examples

### v3

@@ Appendix B White-Box Attacks Supplementary Material
S-R-BP 81.1% S-R-BP 65.1%
S-V-BP 89.2% S-V-BP 64.1%
S-V-T5 90.9% S-V-T5 65.8%
S-V-T10 91.4% S-V-T10 65.4%
S-R-T5 89.2% S-R-T8 59.7%
S-R-T10 91.6% - -
C-101x3 98.7% C-101x3 91.8%
C-V 91.9% C-V 66.6%
C-R 92.1% C-R 61.3%
V-B32 98.6% V-B32 91.7%
V-B16 98.9% V-B16 92.8%
V-L16 99.1% V-L16 94.0%
Table 5: Clean Accuracy for models for CIFAR-10 and CIFAR-100 datasets.
S-V-T5 S-R-BP C-V V-L16 C152x4-512
57.53% 60.82% 71.59% 82.94% 85.31%
Table 6: Clean Accuracy for ImageNet models
Surrogate Gradient
Surrogate gradient (Neftci et al. 2019) has become a popular technique to overcome the non-differentiable problem of spiking neuron’s binary activation. The shapes of the surrogate gradient functions discussed in the paper are shown in Figure 3. Let $u(x)$ be Heaviside step function, and $u^{\prime}(x)$ be its derivative. The surrogate gradients investigated in this work are discussed as follows:
Sigmoid (Bengio et al. 2013) indicates that a hard threshold function’s derivative can be approximated by that of a sigmoid function. The surrogate gradient is given by Equation 12:
$u^{\prime}(x)\approx\frac{e^{\vartheta-x}}{(1+e^{\vartheta-x})^{2}}$ (12)
Erfc (Fang et al. 2020b) proposes to use the Poisson neuron’s spike rate function. The spike rate can be characterized by complementary error function (erfc), and its derivative is calculated as Equation 13, where $\sigma$ controls the sharpness:
Arctan (Fang et al. 2021b) uses gradient of arctangent function as surrogate gradient, which is given by:
$u^{\prime}(x)\approx\frac{1}{1+\pi^{2}(x-\vartheta)^{2}}$ (14)
Piece-wise linear function (PWL) There are various works use PWL function as gradient surrogate (Rathi et al. 2021; Bellec et al. 2018; Neftci et al. 2019). Its formulation is given below:
$u^{\prime}(x)\approx\max(0,1-|x|)$ (15)
Fast sigmoid (Zenke et al. 2018) uses fast sigmoid as a replacement of the sigmoid function, the purpose is to avoid expensive exponential operation and to speed up computation.
$u^{\prime}(x)\approx\frac{1}{1+(1+|x-\vartheta|)^{2}}$ (16)
Piece-wise Exponential (Shrestha et al. 2018) suggests that Probability Density Function (PDF) for a spiking neuron to change its state (fire or not) can approximate the derivative of the spike function. Spike Escape Rate, which is a piece-wise exponential function, can be a good candidate to characterize this probability density. It is given by Equation 17
$u^{\prime}(x)\approx\frac{1}{\alpha e^{-\beta|x-\vartheta|}}$ (17)
where $\alpha$ and $\beta$ are two hyperparamaters.
CIFAR 10
$\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031
PWL 53.3% 26.9% 12.6% 7.0% 3.7%
Erfc 53.7% 26.9% 12.0% 6.4% 3.2%
Sigmoid 95.1% 8.72% 80.1% 73.3% 65.4%
Piecewise Exp. 97.5% 96.0% 93.2% 90.0% 87.9%
Rectangle 55.8% 30.7% 16.4% 9.2% 5.8%
Fast Sigmoid 77.4% 55.3% 37.0% 23.4% 15.3%
Arctan 55.0% 26.4% 11.2% 6.1% 3.1%
CIFAR 100
$\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031
PWL 26.4% 8.4% 3.5% 1.7% 1.0%
Erfc 26.9% 8.4% 3.4% 2.0% 0.9%
Sigmoid 88.9% 77.6% 66.0% 54.0% 44.5%
Piecewise Exp. 92.5% 86.2% 79.5% 7.28% 65.5%
Rectangle 29.7% 10.9% 4.9% 2.5% 1.7%
Fast Sigmoid 63.4% 36.5% 22.5% 13.2% 8.7%
Arctan 28.5% 9.6% 3.8% 1.9% 0.8%
@@
Table 7: White box attack success rate for transfer SNN VGG 16 model on CIFAR 10 and CIFAR 100 with respect to different surrogate gradients.
CIFAR 10
$\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031
PWL 66.5% 37.7% 20.7% 13.9% 9.5%
Erfc 64.1% 37.0% 22.1% 13.1% 7.7%
Sigmoid 85.7% 63.7% 46.3% 30.7% 19.4%
Piecewise Exp. 95.1% 91.0% 84.6% 76.7% 67.5%
Rectangle 76.5% 56.6% 40.6% 30.5% 23.9%
Fast Sigmoid 96.3% 92.5% 87.9% 80.0% 74.1%
Arctan 59.9% 28.9% 13.9% 8.6% 4.6%
CIFAR 100
$\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031
PWL 41.6% 18.6% 8.4% 3.6% 2.1%
Erfc 40.2% 16.0% 7.4% 2.9% 1.5%
Sigmoid 75.0% 58.0% 42.9% 30.9% 19.3%
Piecewise Exp. 82.7% 78.7% 73.2% 64.5% 58.1%
Rectangle 60.4% 38.6% 24.5% 16.7% 11.9%
Fast Sigmoid 86.1% 83.8% 81.5% 78.2% 73.9%
Arctan 34.2% 11.2% 4.6% 2.2% 0.6%
@@
Table 8: White box attack success rate for BP SNN VGG 16 model on CIFAR 10 and CIFAR 100 with respect to different surrogate gradients.
CIFAR 10
$\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031
PWL 80.7% 63.9% 48.3% 34.0% 23.3%
Erfc 76.5% 50.6% 29.6% 15.6% 06.8%
Sigmoid 85.6% 72.2% 56.1% 41.5% 29.0%
Piecewise Exp. 89.6% 84.3% 76.8% 66.8% 57.7%
Rectangle 79.7% 55.7% 37.3% 23.7% 13.1%
Fast Sigmoid 81.5% 61.2% 42.8% 27.6% 16.2%
Arctan 74.4% 49.7% 29.0% 15.6% 7.2%
CIFAR 100
$\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031
PWL 62.3% 32.0% 18.0% 9.9% 4.4%
Erfc 25.8% 5.9% 1.7% 0.6% 0.2%
Sigmoid 68.1% 38.7% 21.8% 12.6% 07.2%
Piecewise Exp. 90.0% 86.9% 79.4% 70.8% 61.9%
Rectangle 32.2% 8.7% 3.4% 0.9% 0.4%
Fast Sigmoid 51.2% 20.7% 10.0% 4.5% 2.4%
Arctan 25.8% 6.3% 1.9% 0.7% 0.2%
@@
PWL 27.5% 5.9 % 1.4% 0.45% 0.1%
Erfc 30.1% 8.6% 2.05% 0.95% 0.25%
Sigmoid 38.85% 10.35% 2.9% 0.6% 0.1%
Piecewise Exp. 18.85% 2.25% 0.25% 0.05% 0.05%
Rectangle 56.25% 29.55% 15.85% 8.55% 5.25%
Fast Sigmoid 32.65% 7.65% 1.3% 0.3% 0.15%
Arctan 20.3% 3.65% 0.35% 0% 0%
@@
Table 9: White box attack success rate for SEW ResNet 18 model on CIFAR 10, CIFAR 100 and ImageNet with respect to different surrogate gradients.
ImageNet
$\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031
PWL 39.65% 13.3% 3.9% 1.4% 0.55%
Erfc 40.65% 15.05% 5.1% 2.05% 0.95%
Sigmoid 46.0% 15.15% 4.35% 0.155% 0.25%
Piecewise Exp. 29.60 % 5.3% 0.9% 0.1% 0%
Rectangle 64.95% 44.8% 30.5% 17.9% 12.15%
Fast Sigmoid 48.8% 17.95% 5.7% 1.7% 0.75%
Arctan 32.2% 9.3% 3.0% 0.85% 0.3%
Table 10: White box attack success rate for Vanilla Spiking ResNet 18 model on ImageNet with respect to different surrogate gradients.
Rectangular function is used by (Wu et al. 2018, 2019) as surrogate gradient. $\alpha$ is a hyperparameter that controls height and width.
White-box Attack Success Rate
We evaluated white-box attack success rate on 3 SNN models (transfer SNN, BP SNN and SEW ResNet) with respect to 7 different surrogate gradients. Results are shown in Table 789 and 10.
Fast Gradient Sign Method
The Fast Gradient Sign Method (FGSM) (Goodfellow et al. 2015) is a white box attack that generates adversarial examples by adding noise to the clean image in the direction of the gradients of the loss function:
$x_{adv}=x+\epsilon\cdot sign(\Delta_{x}L(x,y;\theta))$ (19)
where $\epsilon$ is the attack step size parameter and $L$ is the loss function of the targeted model. The attack performs only a single step of perturbation, and applies noise in the direction of the sign of the gradient of the model’s loss function.
Projected Gradient Descent
The Projected Gradient Descent attack (PGD) (Madry et al. 2018) is a modified version of the FGSM attack that implements multiple attack steps. The attack attempts to find the minimum perturbation, bounded by $\epsilon$, which maximizes the model’s loss for a particular input, $x$. The attack begins by generating a random perturbation on a ball centered at x and with radius $\epsilon$. Adding this noise to $x$ gives the initial adversarial input, $x_{0}$. From here the attack begins an iterative process that runs for $k$ steps. During the $i^{th}$ attack step the perturbed image, $x_{i}$, is updated as follows:
$x_{i}=P(x_{i-1}+\alpha\cdot sign(\Delta_{x}L(x_{i-1},y;\theta))$ (20)
Momentum Iterative Method
The Momentum Iterative Method (MIM) (Dong et al. 2018) applies momentum techniques seen in machine learning training to the domain of adversarial machine learning. Similar to those learning methods, the MIM attack’s momentum allows it to overcome local minima and maxima. The attack’s main formulation is similar to the formulation seen in the PGD attack. Each attack iteration is calculated as follows:
$x_{i}=clip_{x,epsilon}(x_{i-1}+\frac{\epsilon}{t}\cdot sign(g_{i}))$ (21)
where $x_{i}$ represents the adversarial input at iteration $i$, $\epsilon$ is the total attack magnitude, and $t$ is the total number of attack iterations. $g_{i}$ represents the accumulated gradient at step $i$ and is calculated as follows:
$g_{i}=\mu\cdot g_{i-1}+\frac{\Delta_{x}L(x_{i-1},y;\theta)}{||\Delta_{x}L(x_{i-1},y;\theta)||_{1}}$ (22)
where $\mu$ represents a momentum decay factor. Due to its similarity of formulation, the MIM attack degenerates to an iterative form of FGSM as $\mu$ approaches 0.
Auto-PGD
The Auto-PGD (Croce et al. 2020) is a budget-aware step size-free variant of PGD. The algorithm partitions the available $N_{iter}$ iterations to at first search for a good initial point. Then in the exploitation phase, it progressively reduces the step size to maximize the attack results. However, the reduction in step size is not a priori, but is governed by the optimization trend: if the target value grows sufficiently fast, then the step size is most likely appropriate, otherwise it is reasonable to reduce the step size. The gradient update of Auto-PGD follows closely the classic algorithm and only adds a momentum term. Let $\eta_{i}$ be the step size at iteration $i$, then the update step is as follows:
$z_{i+1}=P\left(x_{i}\right.$ $\left.+\eta_{i}\nabla f\left(x_{i}\right)\right)$ (23)
$x_{i+1}=P\left(x_{i}\right.$ $+\alpha\cdot\left(z_{i+1}-x_{i}\right)$
$\left.+(1-\alpha)\cdot\left(x_{i}-x_{i-1}\right)\right)$
where $\alpha\in[0,1]$ regulates the influence of the previous update on the current one.
Figure 4: Accuracy of SEW ResNet on CIFAR 100 with respect to different gradient surrogates.
Figure 5: Accuracy of Vanilla Spiking ResNet on ImageNet with respect to different gradient surrogates.
Dataset CIFAR-10 CIFAR-100 ImageNet
Model S-R-BP S-V-BP S-V-T10 S-V-T5 S-R-BP S-V-BP S-V-T10 S-V-T5 S-R-BP S-V-T5
Timesteps 4 20 10 5 5 30 10 5 4 5
Table 11: The timesteps used for each SNN model we used in the paper
@@
AutoPGD 64.3% 61.7% 63.1% 63.4% 53.0% 64.0% 44.8%
Table 12: White box attack success rate for ResNet-18 SNN model with DM adversarial training method on CIFAR 10, CIFAR 100 with respect to different surrogate gradients.
@@
AutoPGD 90.4% 90.6% 90.6% 90.1% 84.9% 88.7% 82.2%
Table 13: White box attack success rate for ResNet-18 SNN model with FAT adversarial training method on CIFAR 10, CIFAR 100 with respect to different surrogate gradients.
@@
AutoPGD 100.0% 100.0% 100.0% 100.0% 99.8% 99.9% 99.1%
Table 14: White box attack success rate for ResNet-19 SNN model with TIC adversarial training method on CIFAR 10, CIFAR 100 with respect to different surrogate gradients.
CIFAR-10
Arctan PWL Erfc Sigmoid PWE Rectangle Fast Sigmoid STDB
MIM 46.4% 45.9% 45.9% 46.8% 28.5% 49.3% 25.2% 94.8%
PGD 47.9% 45.9% 47.1% 46.8% 27.7% 45.9% 25.6% 96.4%
AutoPGD 73.0% 71.5% 72.3% 73.2% 54.6% 69.7% 55.1% 98.5%
CIFAR-100
Arctan PWL Erfc Sigmoid PWE Rectangle Fast Sigmoid STDB
MIM 69.8% 70.7% 70.4% 71.0% 55.6% 70.5% 50.4% 65.8%
PGD 69.2% 69.6% 71.1% 69.2% 54.3% 69.1% 49.4% 69.1%
AutoPGD 90.4% 90.6% 90.6% 90.1% 84.9% 88.7% 82.2% 68.0%
Table 15: White box attack success rate for VGG-16 SNN on CIFAR 10 and VGG-11 SNN on CIFAR 100 with HIRE adversarial training method with respect to different surrogate gradients.
Appendix C Adversarial Trained SNNs Supplementary Material
To further validate the substantial influence of the surrogate gradient estimator, we consolidate four state-of-the-art adversarial training (AT) methods and conduct training on SNNs to study in our paper. Specifically, we modify two new effective adversarial training methods from CNN domain, namely DM (Wang et al. 2023) and FAT (Zhang et al. 2020a), for SNN-specific training. Additionally, we introduce two newly proposed adversarial training methods for SNNs, denoted as HIRE (Kundu et al. 2021) and TIC (Kim et al. 2023). We adopt the AT for SNNs and perform FGSM, MIM, PGD, and AutoPGD on the trained SNNs with different surrogate gradient estimators.
DM (Wang et al. 2023) proposed to use class-conditional elucidating diffusion model (EDM) (Karras et al. 2022) to generate augmented datasets for CIFAR-10 and CIFAR-100, and use TRADES (Zhang et al. 2019) pipeline for adversarial training. This TRADES training utilizes a classification-calibrated loss theory to attain a differentiable upper bound, enhancing model accuracy while shifting the decision boundary from the data to boost robustness. To train our SEW ResNet18 SNNs on CIFAR-10 and CIFAR-100, we utilized TRADES5 with 10M and 1M augment data as per the original paper settings. Notably, this adversarial training yields the highest robustness among all investigated methods but achieves lowest model accuracy (66.8% for CIFAR-10 and 41.0% for CIFAR-100). The attack results are shown in Table 12.
Friendly adversarial training (FAT) (Zhang et al. 2020a) focuses on identifying adversarial data that minimizes loss. This method trains a DNN by minimizing the loss with wrongly-predicted adversarial data and maximizing the loss with correctly-predicted adversarial data. The implementation employs early-stopped PGD for ease of adaptation. We maintain consistency with the original paper’s approach by employing PGD-10-5 to train SEW ResNet18 SNNs on CIFAR-10 (73.2%) and CIFAR-100 (40.8%). While the trained SNNs demonstrate some level of robustness, it is not as potent as observed in (Wang et al. 2023). The attack results are detailed in Table 13.
Temporal Information Concentration (TIC) (Kim et al. 2023) highlights that in SNNs, the information gradually shifts from later timesteps to earlier timesteps during training. This phenomenon is investigated in this study, where they estimate the Fisher information of the weights and propose a loss function to regulate the trend of Fisher information in SNNs. The adversarial training approximates the relationship between the loss function and Fisher information, pushing the loss value towards a specific target $\alpha$. Following the guidance provided in the paper, we train ResNet19 SNNs with $\alpha=1e-3$ for CIFAR-10 and $\alpha=1e-4$ for CIFAR-100. Although the SNNs achieve high accuracy (92.3% and 72.1%), as stated in the paper, their robustness is not strong, even when different estimators are used, as indicated in Table 14.
HIRE-SNN (Kundu et al. 2021) represents a spike timing dependent backpropagation (STDB) driven SNN training algorithm aimed at enhancing the inherent robustness of SNNs. The approach is tailored to conversion-based SNN training and prioritizes the optimization of model trainable parameters. This is achieved by introducing crafted noise to perturb pixel values across time steps in images. Our implementation follows the original paper’s methodology, dividing time steps into two equal-length intervals and introducing input noise after each period during training. For CIFAR-10 and CIFAR-100 datasets, we trained VGG-16 (89.0%) and VGG-11 (66.1%), respectively. Although the SNNs achieve higher accuracy, they demonstrate very low robustness when the correct surrogate gradient estimator is chosen, as shown in Table 15.
Appendix D Auto-SAGA Supplementary Material
Here we list the attack settings for experiments shown in Section Auto-SAGA Experimental Results. For all attacks, we use a maximum perturbation of $\epsilon=0.031$ with respect to the $l_{\infty}$ norm.

### v4

@@ 3.2 Surrogate Gradient Estimator Experiments
$\epsilon$ $\epsilon$
0.0062 0.0124 0.0186 0.0248 0.031 0.0062 0.0124 0.0186 0.0248 0.031
PWL 26.7% 56.5% 77.3% 83.9% 89.9% 47.6% 80.4% 90.3% 95.7% 98.2%
Erfc 27.6% 60.1% 78.4% 85.6% 90.6% 51.0% 79.2% 91.7% 96.1% 97.7%
Sigmoid 10.6% 21.5% 44.7% 66.8% 78.8% 22.9% 33.7% 46.9% 65.3% 77.4%
Piecewise Exp. 6.1% 6.8% 9.5% 14.7% 19.6% 18.0% 21.4% 22.7% 25.5% 29.5%
Rectangle 17.5% 38.6% 53.3% 68.5% 75.1% 33.3% 51.7% 70.9% 81.1% 86.8%
Fast Sigmoid 4.4% 5.1% 7.9% 13.1% 17.4% 17.8% 16.6% 18.9% 19.8% 19.7%
Arctan 33.3% 67.1% 84.5% 91.3% 94.6% 61.8% 87.4% 95.1% 98.0% 99.0%
@@
Table 3: APGD attack success rate for SEW ResNet 18 model on CIFAR-10, CIFAR-100 and ImageNet with respect to different surrogate gradients.
CIFAR-10 CIFAR-100 ImageNet
$\epsilon$ $\epsilon$ $\epsilon$
Surrogate Grad. 0.0062 0.0124 0.0186 0.0248 0.031 0.0062 0.0124 0.0186 0.0248 0.031 0.0062 0.0124 0.0186 0.0248 0.031
PWL 18.9% 40.6% 67.9% 83.6% 93.0% 72.0% 93.9% 98.2% 99.4% 99.8% 72.5% 94.1% 98.6% 99.6% 99.9%
Erfc 19.9% 43.0% 67.2% 83.7% 92.2% 71.5% 94.4% 98.4% 99.5% 99.8% 69.9% 91.4% 98.0% 99.1% 99.8%
Sigmoid 10.2% 22.6% 33.1% 49.3% 64.3% 25.0% 53.2% 78.0% 86.9% 92.5% 61.2% 89.7% 97.1% 99.4% 99.9%
Piecewise Exp. 16.1% 30.2% 53.8% 72.5% 85.9% 52.8% 85.8% 94.7% 98.2% 99.1% 81.2% 97.8% 99.8% 100.0% 100.0%
Rectangle 17.3% 36.8% 55.1% 74.9% 85.1% 62.8% 89.7% 96.5% 98.5% 99.4% 43.8% 70.5% 84.2% 91.5% 94.8%
Fast Sigmoid 14.4% 28.9% 54.8% 71.6% 83.0% 41.3% 77.9% 90.4% 95.6% 97.9% 67.4% 92.4% 98.7% 99.7% 99.9%
Arctan 20.9% 42.8% 66.8% 83.1% 92.3% 70.8% 94.0% 98.6% 99.4% 99.8% 79.7% 96.4% 99.7% 100.0% 100.0%
@@
Table 4: APGD attack success rate for Vanilla Spiking ResNet 18 model on ImageNet with respect to different surrogate gradients.
@@
PWL 60.4% 86.7% 96.1% 98.6% 99.5%
Erfc 59.4% 85.0% 94.9% 98.0% 99.1%
Sigmoid 54.0% 84.9% 95.7% 98.5% 99.8%
Piecewise Exp. 70.4% 94.7% 99.1% 99.9% 100.0%
Rectangle 35.1% 55.2% 69.5% 82.1% 87.9%
Fast Sigmoid 51.2% 82.1% 94.3% 98.3% 99.3%
Arctan 67.8% 90.7% 97.0% 99.2% 99.7%
Experimental Setup: We evaluate the attack success rate of aforementioned gradient estimators on SNNs trained with and without adversarial training. For the attack, we use one of the most common white-box attacks, the Auto Projected Gradient Descent (Auto-PGD) attack [25] with respect to the $l_{\infty}$ norm. When conducting Auto-PGD, we keep the model’s forward pass unchanged, and the surrogate gradient function is substituted in the backward pass only. For the undefended (vanilla) SNNs we test 3 types of SNNs on CIFAR-10/100 [50] and 2 types of SNNs on ImageNet [51] using 7 different surrogate gradient estimators. We test the Transfer SNN VGG-16 [7], the BP SNN VGG-16 [9], a Spiking Element Wise (SEW) ResNet [10], and Vanilla Spiking ResNet [52].
Vanilla SNN Experimental Analysis: The results of our surrogate gradient estimation experiments are shown in Figure 2. For each model and each gradient estimator, we vary the maximum perturbation bounds from $\epsilon{=}0.0062$ to $\epsilon{=}0.031$ on the x-axis and run the Auto-PGD attack on 1000 (CIFAR-10 and CIFAR-100), and 2000 (ImageNet) clean, correctly identified and class-wise balanced samples from the validation set. The corresponding robust accuracy is then measured on the y-axis. Our results show that unlike what the literature reported for SNN training [35], the choice of surrogate gradient estimator hugely impacts SNN attack performance. In most cases, the arctan yields the lowest accuracy (the highest attack success rate), which includes Transfer SNN C10, BP SNN C10, Transfer SNN C100, BP SNN C100, SEW ResNet C100 experiments.
This trend does not occur for ImageNet, where PWE performs best and arctan performs second best in Vanilla Spiking ResNet and ImageNet experiment. And in SEW ResNet ImafeNet experiment, both PWE and and arctan achieve 100% attack successful rate.
The worst gradient estimator varies in different experiments. For example, though PEW achieves best ASR in two experiment on ImageNet dataset, it has lowest ASR in Transfer SNN C10, BP SNN C10, Transfer SNN C100, BP SNN C100. And sigmoid performs worst in SEW SNN 10 and SEW SNN C100.
Results also show that there is significan performance gap between the best and worst gradient estimator. For example, in BP SNN C10, the arctan achieves 64.6% ASR, while ASR of PWE is merely 19.6%; and in Transfer SNN C100, arctan achieves 99.3% ASR, however PWE only achieves 46.8% ASR.
To reiterate, this set of experiments highlights a significant finding: for SNNs, choosing the right surrogate gradient estimator is critical for achieving a high white-box attack success rate.
3.3 Adversarial Trained SNN Experimental Analysis
To further validate the substantial influence of the surrogate gradient estimator (SG), we consolidate four state-of-the-art adversarial training (AT) methods and conduct training on SNNs in our study. Specifically, we modify two effective adversarial training methods from the CNN domain, namely DM [27] and FAT [14], for SNN training. Additionally, we introduce two newly proposed adversarial training methods for SNNs, denoted as HIRE [19] and TIC [26]. We adopt these AT methods for SNNs and perform MIM, PGD, and Auto-PGD attacks on the trained SNNs using different surrogate gradient estimators. We set the maximum perturbation bounds $\epsilon=0.031$ and attack steps to 40 for all three attacks, with a step size $\epsilon_{\text{step}}=0.01$ for MIM and PGD. We run 1000 clean, correctly identified, and class-wise balanced samples from the validation set on CIFAR-10 and CIFAR-100.
The DM adversarial trained SEW ResNet18 SNNs utilized TRADES5 with 10M and 1M augment data as per the original paper settings on CIFAR-10 and CIFAR-100. Notably, this adversarial training yields the highest robustness among all investigated methods but achieves lowest clean model accuracy (66.8% for CIFAR-10 and 41.0% for CIFAR-100). The attack results are shown in Table 5. For FAT training, the implementation employs early-stopped PGD for ease of adaptation. We maintain consistency with the original paper’s approach by employing PGD-10-5 ($k=10$, and $\tau=5$) to train SEW ResNet18 SNNs on CIFAR-10 (73.2%) and CIFAR-100 (40.8%). While the trained SNNs demonstrate some level of robustness, it is not as robust as results shown for CNNs as presented in [27], especially for Auto-PGD results on CIFAR-100 SNN. This discrepancy could be because the FAT training method is originally designed for CNNs and may not be fully adapted to the SNN settings. The attack results are detailed in Table 6.
@@ 3.3 Adversarial Trained SNN Experimental Analysis
Table 5: White box attack success rate for ResNet-18 SNN model with DM adversarial training method on CIFAR-10, CIFAR-100 with respect to different surrogate gradients.
@@
Auto-PGD 64.3% 61.7% 63.1% 63.4% 53.0% 64.0% 44.8%
Table 6: White box attack success rate for ResNet-18 SNN model with FAT adversarial training method on CIFAR-10, CIFAR-100 with respect to different surrogate gradients.
@@
Auto-PGD 90.4% 90.6% 90.6% 90.1% 84.9% 88.7% 82.2%
To train the SNNs on HIRE SNNs, we follows the original paper’s methodology, dividing time steps into two equal-length intervals and introducing input noise after each period during training. For CIFAR-10 and CIFAR-100 datasets, we trained VGG-16 (89.0%) and VGG-11 (66.1%), respectively. Although the SNNs achieve higher accuracy, they demonstrate very low robustness when the correct surrogate gradient estimator is chosen, as shown in Table 7. As for SNNs with the TIC method, we follow the guidance provided in the paper and train ResNet-19 SNNs with $\alpha=1e-3$ for CIFAR-10 and $\alpha=1e-4$ for CIFAR-100. Although the SNNs achieve high accuracy (92.3% and 72.1%), as stated in the paper, their robustness is not strong, even when different estimators are used, as indicated in Table 8.
Table 7: White box attack success rate for VGG-16 SNN on CIFAR-10 and VGG-11 SNN on CIFAR-100 with HIRE adversarial training method with respect to different surrogate gradients.
CIFAR-10
Arctan PWL Erfc Sigmoid PWE Rectangle Fast Sigmoid STDB
MIM 83.2% 66.8% 67.7% 42.4% 16.9% 47.3% 79.3% 94.8%
PGD 66.7% 49.6% 49.7% 45.3% 17.6% 35.6% 84.9% 96.4%
AutoPGD 91.1% 80.7% 80.7% 65.8% 31.3% 69.1% 93.6% 98.5%
CIFAR-100
Arctan PWL Erfc Sigmoid PWE Rectangle Fast Sigmoid STDB
MIM 93.5% 93.8% 93.5% 29.8% 18.0% 93.0% 74.4% 65.8%
PGD 95.0% 95.1% 94.7% 30.4% 19.4% 94.1% 78.4% 69.1%
AutoPGD 95.4% 95.9% 95.8% 37.9% 26.2% 94.9% 80.8% 68.0%
Table 8: White box attack success rate for ResNet-19 SNN model with TIC adversarial training method on CIFAR-10, CIFAR-100 with respect to different surrogate gradients.
@@
Auto-PGD 100.0% 100.0% 100.0% 100.0% 99.8% 99.9% 99.1%
Figure 3: Attack success rate of Auto-PGD with $\epsilon=0.031$ on adversarially trained SNNs using the best and worst possible Surrogate Gradient Estimator (SG).
Surrogate Estimator Discussion: We summarize the attack success rate using the best and worst possible Surrogate Gradient Estimator (SG) for Auto-PGD with $\epsilon=0.031$ on CIFAR-10/100. It can clearly be seen from Figure. 3 that the choice of estimator is extremely significant in how effective the attack is. If the worst estimator was used, the attack success rate would be significantly lower than if the best estimator was used. For example, for CIFAR-10, for the DM SNN, the difference in attack success rate between the best and worst SG is $21.9\%$. If the worst estimator was used, the attack success rate would be $34.1\%$, whereas if the best estimator was used, the attack success rate is $56.0\%$. Similarly, for HIRE on CIFAR-100, the attack success rate would be $69.7\%$ higher with the best SG compared to the worst SG. If the worst estimator was used, the attack success rate would be $26.2\%$ compared to $95.9\%$ with the best SG. Just like obfuscating gradients gives false robustness [53], improper surrogate gradients can also yield a false sense of security. Our results clearly show the choice of estimator significantly impacts the success of white-box attacks on adversarially trained SNNs, highlighting the need for careful selection of SGs to ensure accurate evaluation of model robustness.
Model Robustness Analysis: In addition to the influence of surrogate gradients, we observe that different adversarial training methods result in significantly varied robustness profiles across SNN models. Specifically, DM and FAT apply defense mechanisms originally designed for CNNs, whereas TIC and HIRE-SNN are tailored explicitly for the temporal and structural dynamics of SNNs. This fundamental difference in design goals and training strategies leads to distinct behaviors under attacks with different surrogate gradients. The DM and FAT-trained models exhibit moderate robustness across a wide range of surrogate gradients, but this general resilience often comes at the cost of lower clean accuracy, as previously discussed. In contrast, the SNN-specific adversarial training methods, TIC and HIRE, demonstrate more pronounced differences. HIRE-SNN, in particular, achieves strong robustness against certain surrogate gradients, yet it is highly vulnerable when attacked with carefully chosen, effective SGs. This indicates that while HIRE can deliver higher utility and robust performance in specific scenarios, its overall robustness is inconsistent. TIC-trained models, on the other hand, show uniform vulnerability across surrogate gradients, suggesting that the temporal concentration of Fisher information alone is insufficient to withstand adaptive attacks. These concrete differences in robustness motivate further investigation into transferability among original and adversarially trained SNNs, and underscore the importance of developing adaptive and generalizable attack methods such as our proposed Mixed Dynamic Spiking Estimation (MDSE) Attack.
4 SNN Transferability Study
In this section, we investigate two fundamental security questions pertaining to SNNs: